OVHCloud Web Hosting Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
FS#9416 — HAU59
Incident Report for Web Cloud
Resolved
We have a problem on the hau59 optic loop.
The DSLAM is not accessible at the moment.
The issue is currently being analysed.

The hau59 lines are synchronised but there is no connection.

Update(s):

Date: 2013-09-27 14:52:38 UTC
A splicing issue has just been corrected.

The fibre is up again and the MDT is operational.

Date: 2013-09-27 14:04:46 UTC
The cut has been detected by the fibre operator. The technicians are accessing the optic chamber.

Date: 2013-09-27 13:03:03 UTC
The teams are still investigating.

Date: 2013-09-27 10:38:54 UTC
We found a cut on the 7km fibre, 3km from the MDF.
The fibre maintenance team will arrive at 13:30.

Date: 2013-09-27 10:09:47 UTC
A team in currently on site.
Posted Sep 27, 2013 - 07:57 UTC